OpenStudy (anonymous):

To cut one shelf 24 centimetres long and two shelves 65 centimetres long from a single board, how long must the board be?

OpenStudy (ganpat):

look like this way.. big = small + small.. SO u want to find how big ?? Big = 24 + 2 (65) = 24 + 130 = 154 cm...